OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (13, -3) and (10, 12). Please show all of your work.

OpenStudy (anonymous):

find the slope m = (12 + 3)/(10 - 13) = -5 Lines can be written as y = mx + b and in this case y = -5x + b or 12 = -50 + b or b = 62

OpenStudy (anonymous):

ty